# DiffQRCoder: Diffusion-based Aesthetic QR Code Generation with Scanning Robustness Guided Iterative Refinement

### v1

@@ Abstract
With the success of Diffusion Models for image generation, the technologies also have revolutionized the aesthetic Quick Response (QR) code generation. Despite significant improvements in visual attractiveness for the beautified codes, their scannabilities are usually sacrificed and thus hinder their practical uses in real-world scenarios. To address this issue, we propose a novel Diffusion-based QR Code generator (DiffQRCoder) to effectively craft both scannable and visually pleasing QR codes. The proposed approach introduces Scanning-Robust Perceptual Guidance (SRPG), a new diffusion guidance for Diffusion Models to guarantee the generated aesthetic codes to obey the ground-truth QR codes while maintaining their attractiveness during the denoising process. Additionally, we present another post-processing technique, Scanning Robust Manifold Projected Gradient Descent (SR-MPGD), to further enhance their scanning robustness through iterative latent space optimization. With extensive experiments, the results demonstrate that our approach not only outperforms other compared methods in Scanning Success Rate (SSR) with better or comparable CLIP aesthetic score (CLIP-aes.) but also significantly improves the SSR of the ControlNet-only approach from 60% to 99%. The subjective evaluation indicates that our approach achieves promising visual attractiveness to users as well. Finally, even with different scanning angles and the most rigorous error tolerance settings, our approach robustly achieves over 95% SSR, demonstrating its capability for real-world applications.
@@ 1 Introduction
To address the instability of scanability in previous generative-based methods, we propose Diffusion-based QR Code generator (DiffQRCoder), a diffusion-based approach to balance the scannability and aesthetics of QR codes. We introduce Scanning Robust Loss (SRL), specifically designed for evaluating the scannability of a beautified QR code with respect to its reference code. Building on SRL, we develop Scanning Robust Perceptual Guidance (SRPG), an extension of the Classifier Guidance concept [6, 44, 2], which ensures generation fidelity to ground-truth QR codes while preserving aesthetics during the denoising process.
@@ 2.2.2 Generative-based Models
With the rise of diffusion-based image manipulation and conditional control techniques, previous works such as QR Diffusion [15], QR Code AI Art [21] and QRBTF [18] have leveraged generative power of diffusion models to create aesthetic QR codes, primarily relying on ControlNet [45] for guidance. However, more fine-grained guidance that aligns with the inherent mechanisms of QR codes remains unexplored. Another non-open-source method Text2QR [40], introduced a three-stage pipeline that first generates an unscannable QR code using ControlNet, followed by an optimization technique independent of the diffusion process to enhance scannability.
In contrast, our objective is to design training-free, fine-grained guidance that enhances the diffusion denoising process and integrates scannability from the start of generation.
@@ 3.1 Scanning Robust Loss
Given a normalized image $\mathbf{x}$ [^1] Each color channel of an image is divided by 255 to normalize pixel values to the range of 0 to 1, a target QR code $\mathbf{y}$, and a grayscale conversion operator $\mathcal{G}$ (cf., Appendix A). We formulate a pixel-wise error matrix $\mathbf{E}$ that calculates the differences in pixel values between $\mathbf{y}$ and $\mathcal{G}(\mathbf{x})$. $\mathbf{E}$ is calculated as follows:
@@
ZXing [37], a popular barcode scanning library, notes that only the central pixels of each module are essential for decoding QR codes. According to Chu et al. [3], each module is divided into $3\times 3$ submodules. They observed that a QR code remains scannable if the binarized average pixel value in the central submodule matches the center pixel value of the target module [^2] A threshold of $0.5$ is set for binarization. This observation enables the creation of visual variations in the peripheral submodules.
@@ 3.2.2 Stage-2
In stage-2 we adopt the unscannable QR code $\hat{\mathbf{x}}$ generated in stage-1 as a starting point for enhancing scannability and a regularizing reference for LPIPS to preserve aesthetics. First, we convert image $\hat{\mathbf{x}}$ into a latent representation $\tilde{\mathbf{z}}_{t}$ using the VAE encoder and by adding Gaussian noise. We also transform the target QR code $\mathbf{y}$ to be a more similar pattern as $\hat{\textbf{x}}$  [^3] The aesthetic QR code generated by Qart is a binary image, so we can utilize SRL to evaluate the discrepancy between this code and the one we generated. for better conditioning. Both $\tilde{\mathbf{z}}_{t}$ and $\tilde{\mathbf{y}}$ are then fed into ControlNet. The predicted clean latent at each timestep $t$ can be calculated as
@@ 3.3 Post-processing with Scanning-Robust Manifold Projected Gradient Descent (SR-MPGD)
SR-MPGD is a post-processing technique proposed to enhance scanning robustness further. Our goal is to minimize $\mathcal{L}_{\text{SR}}(\mathbf{x},\mathbf{y})$ while ensuring the refined QR code $\mathbf{x}$ still lies on nature image manifold $\mathcal{M}$. The optimization problem is defined as:
$\min_{\mathbf{x}\in\mathcal{M}}\mathcal{L}_{\text{SR}}(\mathbf{x},\mathbf{y}).$ (15)
This constrained optimization problem can be solved via the Projected Gradient Descent (PGD) algorithm. Inspired by the manifold-preserving nature proposed in [10], we use the pre-trained VAE encoder $\mathcal{E}(\cdot)$ to project the image to its space and then iteratively refine the latent by:
@@ 4.1 Experimental Settings
In our experiments, 100 text prompts are generated by GPT-4 as the conditions for Stable Diffusion, consistently using easynegative as the negative prompt. We employ Cetus-Mix Whalefall  [^4] https://huggingface.co/fp16-guy/Cetus-Mix_Whalefall_fp16_cleaned as the checkpoint for Stable Diffusion [30] and QR code Monster v2 [24] as the checkpoint for ControlNet [46], with the guidance_scale set to 1.35 for the latter. We compare with QR code AI Art [21], QR Diffusion  [15], and QRBTF  [18]. The reason we don’t compare with non-generative-based methods is that they rely on reference aesthetic images as input, here we only have prompts describing the aesthetic scenes. And other previous works are unavailable for fair comparison due to non-open-source. Compared models are accessed through their web API using their recommended settings. Detailed parameter settings are provided in Appendix D.
In our QR code setups, we use Version 3 QR codes configured with a medium (M) error correction level and mask pattern 4. Each code includes an 80-pixel padding, and each module is in the size of $20\times 20$ pixels. Additionally, the text message Thanks reviewer! is encoded into our QR codes for most experiments in the paper. We also conduct quantitative and qualitative results in different messages for QR code generation.
We conduct our experiments using a single NVIDIA RTX 4090 GPU. Generating aesthetic QR codes takes approximately 14 to 18 seconds in our two-stage pipeline, each with 40 inference steps.
@@ 5 Conclusion
In this paper, we introduce a novel diffusion-based QR code generator (DiffQRCoder). We propose Scanning-Robust Loss (SRL) to enhance the scannability of QR codes and have established a bridge between SRL and our proposed Scanning-Robust Perceptual Guidance (SRPG). We develop a two-stage generation pipeline with iterative refinement that integrates our SRPG to generate aesthetic QR codes. Moreover, we introduce Scanning-Robust Manifold Projected Gradient Descent (SR-MPGD) to ensure scannability further. Compared to other existing methods, our approach significantly enhances the SSR of the generated codes without compromising visual attractiveness. Thus, the aesthetic QR codes generated through our approach are capable of applications in real-world scenarios.

### v2

@@ Abstract
With the success of Diffusion Models for image generation, the technologies also have revolutionized the aesthetic Quick Response (QR) code generation. Despite significant improvements in visual attractiveness for the beautified codes, their scannabilities are usually sacrificed and thus hinder their practical uses in real-world scenarios. To address this issue, we propose a novel training-free Diffusion-based QR Code generator (DiffQRCoder) to effectively craft both scannable and visually pleasing QR codes. The proposed approach introduces Scanning-Robust Perceptual Guidance (SRPG), a new diffusion guidance for Diffusion Models to guarantee the generated aesthetic codes to obey the ground-truth QR codes while maintaining their attractiveness during the denoising process. Additionally, we present another post-processing technique, Scanning Robust Manifold Projected Gradient Descent (SR-MPGD), to further enhance their scanning robustness through iterative latent space optimization. With extensive experiments, the results demonstrate that our approach not only outperforms other compared methods in Scanning Success Rate (SSR) with better or comparable CLIP aesthetic score (CLIP-aes.) but also significantly improves the SSR of the ControlNet-only approach from 60% to 99%. The subjective evaluation indicates that our approach achieves promising visual attractiveness to users as well. Finally, even with different scanning angles and the most rigorous error tolerance settings, our approach robustly achieves over 95% SSR, demonstrating its capability for real-world applications. Our project page is available at https://jwliao1209.github.io/DiffQRCoder.
@@ 1 Introduction
To address the instability of scanability in previous generative-based methods, we propose Diffusion-based QR Code generator (DiffQRCoder), a training-free approach to balance the scannability and aesthetics of QR codes. We introduce Scanning Robust Loss (SRL), specifically designed for evaluating the scannability of a beautified QR code with respect to its reference code. Building on SRL, we develop Scanning Robust Perceptual Guidance (SRPG), an extension of the Classifier Guidance concept [6, 44, 2], which ensures generation fidelity to ground-truth QR codes while preserving aesthetics during the denoising process.
@@ 2.2.2 Generative-based Models
With the rise of diffusion-based image manipulation and conditional control techniques, previous works such as QR Diffusion [15], QR Code AI Art [21] and QRBTF [18] have leveraged generative power of diffusion models to create aesthetic QR codes, primarily relying on ControlNet [45] for guidance. However, more fine-grained guidance that aligns with the inherent mechanisms of QR codes remains unexplored. Another non-open-source method Text2QR [40], introduced a three-stage pipeline that first generates an unscannable QR code using pre-trained ControlNet, followed by their proposed SELR independent of the diffusion sampling process to ensure scannability. However, the diffusion sampling process of Text2QR can not guarantee scannability on its own; our approach aims to fill this gap by designing a training-free, fine-grained guidance integrating scannability criteria into the diffusion sampling process.
@@ 3.1 Scanning Robust Loss
Given a normalized image $\mathbf{x}$, a target QR code $\mathbf{y}$, and a grayscale conversion operator $\mathcal{G}$ (cf., Appendix A). We formulate a pixel-wise error matrix $\mathbf{E}$ that calculates the differences in pixel values between $\mathbf{y}$ and $\mathcal{G}(\mathbf{x})$. $\mathbf{E}$ is calculated as follows:
@@
ZXing [37], a popular barcode scanning library, notes that only the central pixels of each module are essential for decoding QR codes. According to Chu et al. [3], each module is divided into $3\times 3$ submodules. They observed that a QR code remains scannable if the binarized average pixel value in the central submodule matches the center pixel value of the target module.This observation enables the creation of visual variations in the peripheral submodules.
@@ 3.2.2 Stage-2
In stage-2 we adopt the unscannable QR code $\hat{\mathbf{x}}$ generated in stage-1 as a starting point for enhancing scannability and a regularizing reference for LPIPS to preserve aesthetics. First, we convert image $\hat{\mathbf{x}}$ into a latent representation $\tilde{\mathbf{z}}_{t}$ using the VAE encoder and by adding Gaussian noise. We also transform the target QR code $\mathbf{y}$ to be a more similar pattern as $\hat{\textbf{x}}$ for better conditioning. Both $\tilde{\mathbf{z}}_{t}$ and $\tilde{\mathbf{y}}$ are then fed into ControlNet. The predicted clean latent at each timestep $t$ can be calculated as
@@ 3.3 Post-processing with Scanning-Robust Manifold Projected Gradient Descent (SR-MPGD)
SR-MPGD is a post-processing technique proposed to enhance scanning robustness further. Our goal is to minimize $\mathcal{L}_{\text{SR}}(\mathbf{x},\mathbf{y})$ while ensuring the refined QR code $\mathbf{x}$ still lies on nature image manifold $\mathcal{M}$. The optimization problem is defined as: $\min_{\mathbf{x}\in\mathcal{M}}\mathcal{L}_{\text{SR}}(\mathbf{x},\mathbf{y})$. This constrained optimization problem can be solved via the Projected Gradient Descent (PGD) algorithm. Inspired by the manifold-preserving nature proposed in [10], we use the pre-trained VAE encoder $\mathcal{E}(\cdot)$ to project the image to its space and then iteratively refine the latent by:
@@ 4.1 Experimental Settings
In our experiments, 100 text prompts are generated by GPT-4 as the conditions for Stable Diffusion, consistently using easynegative as the negative prompt. We employ Cetus-Mix Whalefall as the checkpoint for Stable Diffusion [30] and QR code Monster v2 [24] as the checkpoint for ControlNet [46], with the guidance_scale set to 1.35 for the latter. We compare with QR code AI Art [21], QR Diffusion  [15], and QRBTF  [18]. The reason we don’t compare with non-generative-based methods is that they rely on reference aesthetic images as input, here we only have prompts describing the aesthetic scenes. And other previous works are unavailable for fair comparison due to non-open-source. Compared models are accessed through their web API using their recommended settings. Detailed parameter settings are provided in Appendix D.
In our QR code setups, we use Version 3 QR codes configured with a medium (M) error correction level and mask pattern 4. Each code includes an 80-pixel padding, and each module is in the size of $20\times 20$ pixels. Additionally, the text message Thanks reviewer! is encoded into our QR codes for most experiments in the paper. We also conduct quantitative and qualitative results in different messages for QR code generation. We conduct our experiments using a single NVIDIA RTX 4090 GPU. Generating aesthetic QR codes takes approximately 14 to 18 seconds in our two-stage pipeline, each with 40 inference steps.
@@ 5 Conclusion
In this paper, we introduce a novel training-free Diffusion-based QR Code generator (DiffQRCoder). We propose Scanning-Robust Loss (SRL) to enhance QR code scannability and establish its connection with Scanning-Robust Perceptual Guidance (SRPG). Our two-stage generation pipeline with iterative refinement integrates SRPG to produce aesthetic QR codes. Additionally, we introduce Scanning-Robust Manifold Projected Gradient Descent (SR-MPGD) to further ensure scannability. Compared to existing methods, our approach significantly improves SSR without compromising visual appeal and is competent for real-world applications.